The equation for photon energy, E, is
E=hcλ
where h = 6.626×10−34 J⋅s (Planck's constant) and c = 2.99×108 m/s (the speed of light).
What is the wavelength, λ, of a photon that has an energy of E = 3.37×10−19 J ?
1
Expert's answer
2017-08-30T09:52:07-0400
First we have to fix a mistake in the given equation. The correct expression for photon energy looks like this:

E=hc/λ (1)

Express wavelength from (1):

λ = hc/E
Do the calculation:

λ = hc/E = (6.626*10-34 Js * 2.99*108 m/s) / 3.37*10-19 J = 5.88*10-7 m.

This is yellow light.

Answer:
The wavelength of a photon is 5.88*10-7 m (588 nm, yellow light).